# Ionic association and Wien effect in 2D confined electrolytes

### v1

@@ I.4 Ionic current
where $v_{+}$ and $v_{-}$ are respectively the velocity of the cations and the anions along the direction of the field. The velocities are computed every $10^{3}$ time-steps. We estimate the uncertainty of results by splitting the whole trajectories into 10 windows, and by computing the standard deviation of the window averages.
Upon integration across the slit, the ionic current intensity is given by:
$I=hwj.$ (17)
For non interacting ions, the current is given by:
$j_{0}=\frac{2N}{h\mathcal{A}}\frac{(Ze)^{2}D}{k_{B}T}E,$ (18)
which is the regular Ohm’s law. We finally compute the conductivity using:
$\sigma(E)=\frac{j(E)}{E},$ (19)
and we define the Ohmic conductivity:
$\sigma_{0}=\frac{2N}{h\mathcal{A}}\frac{(Ze)^{2}D}{k_{B}T}$ (20)
In Brownian dynamics simulations, as we have chosen the ionic mobility and the ionic charges to be constant across all the simulations, the Ohmic current and conductivity are the same for all $T^{*}$.
I.5 Ionic clusters
When the correlations between the ions becomes strong enough, some ions can be considered as bonded. Then, the distribution of the ions inside the system can be seen as node in a graph, in which bonded ions corresponds to connected nodes. We can define ionic cluster as the connected components of this graph, using standards tools from graph theory Hagberg, Swart, and Schult (2008). This is particularly relevant in MD simulations.
We define $N_{l}$ as the number of clusters of size $l$, and $n_{l}$ the density of ions in clusters of size $l$:
$n_{l}=\frac{lN_{l}}{2N}.$ (21)
The conservation of the total number of ions yields:
$\sum_{l}n_{l}=\frac{1}{2N}\sum_{l}lN_{l}=1.$ (22)
If the number of cations and anions is different inside a given cluster, the cluster will have a net charge $z_{c}e$. It is in particular always the case for clusters with an odd number of ions. We will define the free carrier density as:
$n_{f}^{\text{cluster}}=\frac{1}{2N}\sum_{c}\frac{z_{c}^{2}}{Z^{2}}.$ (23)
In practice, we only observe clusters with charges 0 or ±$\pm Z$. We hypothesise that clusters with multiple charge defects are too short-lived to be relevant. In this case the above formula can be simplified:
$n_{f}^{\text{cluster}}\sim\frac{N_{c}}{2N}=\frac{\sum_{l\text{ odd}}N_{l}}{2N},$ (24)
where $N_{c}$ is the total number of cluster.
II Ion pairing and screening
@@
In this section, we study the formation of ionic pairs using both Brownian dynamics simulations and theoretical analysis. In particular, we determine the average number of pairs both at thermal equilibrium, but also in non-equilibrium conditions under an external field. We show that the system undergoes a conductor/insulator phase transition reminiscent of the 2D Kosterlitz-Thouless transition. We analytically predict the system’s critical temperature, which is found to slightly deviate from the usual KT result due to the quasi-2D nature of confined electrolytes.
II.1 Bjerrum pairing
In both MD and BD simulations, we observed that oppositely charged ions often forms tightly bound pairs (see Figure 1.(b)). These pairs are absent in bulk aqueous systems, and are therefore a direct consequence of nanoscale confinement.
@@ II.1 Bjerrum pairing
The strength of the ionic interactions can be quantified by the Bjerrum lengthBjerrum (1926) $l_{B}$, defined as:
$\beta V(l_{B})=k_{B}T.$ (25)
It is the typical length which compares ion interaction with thermal energy. For a 3D system, the Bjerrum length is defined as
$l^{\text{bulk}}_{B}=\frac{e^{2}}{4\pi\epsilon_{0}\epsilon_{w}k_{B}T},$ (26)
while for 2D systems, a rough estimate is
$l^{\text{2D}}_{B}\sim\frac{\xi}{T^{*}}.$ (27)
@@
$\frac{l^{\text{2D}}_{B}}{l^{\text{bulk}}_{B}}\simeq\frac{\varepsilon_{w}}{\varepsilon_{s}}\gg 1.$ (28)
In bulk water, $l_{B}^{\text{bulk}}=0.7$nm, meaning that long-range interactions between ions are essentially negligible compared to thermal noise. In the slit this is not the case anymore, as the Bjerrum length is orders of magnitude larger. The possibility for the ions to form pairs, called Bjerrum pairs, has to be taken into account in our analysis.
In order to quantify ionic pairing, we can introduce the free ion fraction $n_{f}(E)$ and the ion pair fraction:
$n_{p}(E)=1-n_{f}(E).$ (29)
$n_{f}=0$ and $n_{f}=1$ correspond respectively to fully paired and fully dissociated electrolytes. This fraction can be separated into 2 contributions: a proportion $n_{f}(0)$ of ions which remain free, even in the absence of external field, and a quantity $\delta n_{f}(E)$ coming from additional pairs breaking under the effect of the electric field, in a process known as the second Wien effectKaiser (2014); Onsager (1934):
$\delta n_{f}(E)=n_{f}(E)-n_{f}(0).$ (30)
In what follows, we study the properties of both $n_{f}(0)$ and $\delta n_{f}(E)$. The former is related to the system’s conductivity at vanishingly low electric field, while the latter describes non-linear conduction outside of equilibrium.
II.2 Determination of the free ion fraction at equilibrium
II.2.1 Definition of the charge carrier density
Pairs being neutral, they do not contribute to conduction under an external field. By applying a constant electric field in Brownian simulations, we can obtain the fraction of charge carriers by comparing the measured ionic current to the Ohmic current:
$n_{f}(E)=\frac{j(E)}{j_{0}(E)}.$ (31)
This definition of an ionic pair has the advantage of being the closest to what could be measured in experiments, in addition to being very convenient to compute. In fact, it effectively counts the fraction of charge carriers only if we can neglect higher-order correlations between ions (which tend to impede conduction compared to the Ohmic case), or equivalently if we assume the mobility of individual ions not to vary with the field. We make this assumption in the following, since we expect the non linearity of the system to be dominated by the variation of the fraction of charge carriers (second Wien effect), and not by the variation of the mobility (first Wien effectKaiser (2014); Kaiser et al. (2013), and Debye-Hückel correlations in generalBerthoumieux, Démery, and Maggs (2024)).
However, Equation (31) is not suited for the measure of the ionic current under a very weak electric field, when the signal-to-noise ratio becomes too low. We attribute this noise to thermal fluctuations, as well as finite-size effects, which become particularly prominent when the proportion of free ions is small (i.e. at low temperature and low applied field). In general, this formula gives an accurate result when $E\gtrsim 2\times 10^{7}$ V/m, which is unfortunately too large for a direct estimate of $n_{f}(0)$.
In the next paragraph, we develop an alternative method to estimate the value of the free ion fraction in the absence of external field. We will make use of an alternate definition of ion pairs, based on correlation functions. We show that this definition is robust regardless of the magnitude of the external field. We will therefore use it as a proxy to estimate the free ion fraction at equilibrium.
II.2.2 Determination from correlation functions
We define a typical length scale $d(E)$ associated with the formation of pairs through the correlation function. We assume a separation of the ions into 2 different types of cluster, the free ions and pairs. By splitting the sum in the definition of $g_{\text{op}}(r)$ between paired and unpaired ions, we can rewrite:
$N_{\text{op}}(r,E)=n_{p}\int_{0}^{r}\rho_{\text{intra}}(r)\text{d}r+(N-n_{p})\int_{0}^{r}\rho_{\text{inter}}(r)\text{d}r,$ (32)
where $\rho_{\text{intra}}$ and $\rho_{\text{inter}}$ are respectively the correlation functions of ions within a same cluster (short distances) and correlations of ions in different clusters (large distances). Both of them vary with the external field. When the overlap between those distributions is small, we can define an intermediate length $d(E)$ between intra and inter cluster correlation, which corresponds to the maximal size of a pair. When $r\simeq d(E)$, the first integral is close to 1, when the second one remains close to 0. This yields:
$N_{\text{op}}(d(E),E)=n_{p}(E).$ (33)
In other words, $N_{\text{op}}$ exhibits a plateau (see Figure S2), associated to a minimum of $g_{\text{op}}$. For $T^{*}\gtrsim 0.1$, we do not observe any minima in the correlation functions, which indicates that a geometric criterion is not adapted for our system. Yet, as the fraction of charge carrier is still well defined in presence of external field, we can use this number to define a typical distance between paired ions $d(E)$:
$N_{\text{op}}(d(E>0),E>0)=1-n_{f}(E>0),$ (34)
where $n_{f}(E>0)$ is computed from transport simulations using equation (31). We then extrapolate the value of the pair distance without external field $d(E=0)$, by assuming that the pair distance decays exponentially with the field. We can finally estimate the free ion fraction without external field as the value of the running coordination function at $d(0)$. The whole procedure is shown for $T^{*}=0.4$ in Figure S1.
II.3 Pairing transition
II.3.1 Comparison with the 2D Coulomb gas
We now use the method described above to compute the free ion fraction without an external field as a function of temperature $T^{*}$. Below $T^{*}\lesssim 0.15$, there is a large uncertainty as the measured currents are small. However, in this regime, the approach using the coordination function demonstrate the absence of any free ion, see Figure S2): the system is fully paired.
@@ II.3.1 Comparison with the 2D Coulomb gas
Figure 2: Pairing transition and evolution of the free ion fraction with the electric field (a) (Red points) Free ion fraction from BD simulation for various reduced Coulomb temperature. (Red lines) Fit of the free ion fraction for an infinite order phase transition. (Blue points and line) Debye length obtained from the the free ion fraction using Equation (35). (b) (Circle) Free ion fraction as a function of the electric field, obtained from the measure of the ionic current, at $T^{*}=0.1$. The main source of uncertainty is the uncertainty on the current. (Diamond) Value at equilibrium. (c) Schematic of the phase transition. At small reduced Coulomb temperature or equivalently high ionic interaction, the system is fully paired. Otherwise, there is some remaining free ions. (d) (Light blue circle) Free ion fraction as a function of the electric field, obtained from the measure of the ionic current, at $T^{*}=0.4$. The uncertainty is small as we measure a larger current. (Light blue diamond) Value at equilibrium, see (a). (Blue circle) Contribution from the non equilibrium free ion only, obtained by subtracting the value at equilibrium to the total current. Its uncertainty is dominated by the uncertainty of the equilibrium value.
In order to better understand the behaviour of our system, it is interesting to compare to the case of a purely 2D Coulomb gas $\xi\to\infty$. In this ideal system, there is a phase transition between a fully paired system and a mixture of free ions and pairs. This transition is identical to the Kosterlitz-Thouless transition Kosterlitz and Thouless (1973); Levin (2002) for the 2D XY model, with a same critical temperature $T^{*}=0.25$. A full analytical treatment of this transition has been detailed elsewhere Levin (2002); in the following, we only sketch the main arguments behind this comparison.
The free energy cost of breaking a pair at thermal equilibrium is roughly the electrostatic free energy of creating the Debye correlation cloud around each ion forming the pair. Indeed, every free ion is surrounded by an atmosphere of opposite charge, extending over a typical scale known as the Debye length:
$\lambda_{\text{D}}=\sqrt{\frac{T^{*}}{4\pi n_{f}\rho}}.$ (35)
In the limit where the ionic concentration is not too highRotenberg, Bernard, and Hansen (2018), it should corresponds to the correlation length of electrostatic interactions in the system i.e. the decay length of the correlation function between free ions only.
Overall, the free energy of breaking a pair scales like:
$\beta\Delta F\sim-\frac{1}{T^{*}}\log\lambda_{\text{D}}/r_{0},$ (36)
where $r_{0}$ is the ionic size (here set by the short distance cut-off $C$ of the interaction potential). The chemical equilibrium between pairs and free ions then read:
$n_{p}=1-n_{f}=Kn_{f}^{2}e^{-\beta\Delta F}\propto n_{f}^{2-1/2T^{*}},$ (37)
with $K$ a constant that depends on the exact criterion used to define pairs (e.g., on the threshold in distance under which two ions are assumed to form a pair). For $T^{*}<1/4$, this equation cannot be solved for $n_{f}$ since $02-1/2T^{*}<$ 0. This indicates that chemical equilibrium between pairs and free ions no longer holds, and the system undergoes a phase transition where all ions pair up.
Figure 2 shows the evolution of the free ion fraction with the temperature obtained in our simulations (red symbols). We observe that the free ion fraction decreases when we increase the interaction between the ions, and vanishes for $T_{c}^{*}\simeq 0.17$, a value rather close to the theoretical prediction $T^{*}_{c}=1/4$. We will show in next section that this slight difference can be attributed to deviations to the perfect Coulomb gas (limit $\xi\to\infty$) in our BD simulations ($\xi=13.8$ nm)
Let us now characterize the nature of this phase transition. A hallmark of the Kosterlitz-Thouless transition is that it is of infinite order – in other words, it does not entail a symmetry breaking, and the system’s correlation length depends in a non-algebraic way on the temperature above the transition. Again, this can be shown through simple arguments as follows. For $T^{*}$ sligthly above $T^{*}_{c}$, the system is still almost fully paired, so that $n_{p}\simeq 1$. From the chemical equilibrium (Equation (37)), we obtain that:
$n_{f}\sim K_{2}^{-1/(2-1/2T^{*})}\sim e^{-\frac{T^{*}/2}{T^{*}-T^{*}_{c}}},$ (38)
Where $K_{2}$ is a temperature-independent constant (which depends on $K$). Since $\lambda_{\text{D}}\sim n_{f}^{-1/2}$, we have, for $T^{*}\gtrsim T^{*}_{c}$:
$\lambda_{\text{D}}\sim e^{\frac{T^{*}/4}{T^{*}-T^{*}_{c}}},$ (39)
which shows that indeed the correlation length is non-algebraic (in particular, no critical exponent can be defined).
We represent the value of $\lambda_{\text{D}}$ as computed in the simulation in Figure 2 (blue symbols). We fit its temperature dependence above the transition with the ansatz:
$\lambda_{D}(T^{*}\to T_{c}^{*+})\sim Ae^{\frac{T^{*}/4}{T^{*}-T_{c}^{*}}},$ (40)
where $A$ and $T_{c}^{*}$ are chosen to reproduce our data. The red and blue curves in Figure 2 shows the result of this model. We obtain a very good agreement for a critical temperature $T_{c}^{*}\sim 0.17$. In particular, we recover the non-algebraic nature of the correlation length.
Despite this difference in the exact value of the critical temperature, we can conclude that the phase transition observed in the simulations is indeed in the same universality class as the one of the exact 2D Coulomb gas, or of the Kosterlitz-Thouless transition. In particular, we find that it is a transition of infinite order, as shown by the non-algebraic decay of the correlation length.
II.3.2 Theoretical analysis with finite dielectric length
In the case where confinement is imperfect, $\xi<\infty$ and the above transition is found to occur at a lower temperature than predicted by the perfect Coulomb gas ($\xi\to\infty$). In this section, we provide a theoretical analysis of this phenomenon and compute the approximate value of the transition temperature. Assuming that ions have a radius $r_{0}$ and cannot interpenetrate, we find that at the Debye-Hückel level, the ionic atmosphere surrounding each ions creates an electrostatic potential $\psi$ on ions:
$\psi=\frac{1}{T^{*}}\left[\frac{\xi}{\kappa_{\text{D}}r_{0}(r_{0}+\xi)}\frac{K_{0}(\kappa_{\text{D}}r_{0})-K_{0}(\kappa_{\text{D}}(r_{0}+\xi))}{K_{1}(\kappa_{\text{D}}r_{0})-K_{1}(\kappa_{\text{D}}(r_{0}+\xi))}\right],$ (41)
with $\kappa_{\text{D}}$ the inverse Debye length. Note that in the case of the perfect 2D Coulomb gas, the same quantity reads:
$\psi=\frac{1}{T^{*}}\left[\frac{K_{0}(\kappa_{\text{D}}r_{0})}{\kappa_{\text{D}}r_{0}K_{1}(\kappa_{\text{D}}r_{0})}\right].$ (42)
By using a Debye charging process, one can show that the excess chemical potential of free ions is directly linked to $\psi$:
$\mu_{e}=\frac{1}{2}\psi.$ (43)
The number of free ions is then given by, assuming pairs behave like ideal solutes:
$n_{p}=1-n_{f}=Kn_{f}^{2}e^{2\mu_{e}}.$ (44)
In the case $\xi<\infty$, the solution of Equation (44) cannot be obtained analytically, but can be solved numerically. In Figure 3.(a) we plot the evolution of the free ion fraction with the density above ($T^{*}=0.15$) and below ($T^{*}=0.35$) the critical temperature.
At low temperature, we observe a discontinuity in the free ion fraction at a density $\rho_{t}(T^{*})$ (see red curve in 3). This defines a transition line between a system with a low fraction of free ion below the line, and a large fraction of free ion above. This behaviour is very similar to what is obtained in the perfect 2D case. The main difference is that instead of a fully paired phase $n_{f}=0$, few ions can remain dissociated.
In the perfect 2D case, it has been shownLevin (2002) that the transition line stops with a tricritical point $(T_{c}^{*},\rho_{t}(T_{c}^{*}))$ at the critical temperature $T_{c}^{*}=0.25$. Indeed, when $\rho<\rho_{t}(T_{c}^{*})$ and $T^{*}<T_{c}^{*}$, Equation (44) has no solution: The system is fully paired. This defines a critical line at $T^{*}=T_{c}^{*}$ that stops a the tricritical point. By analogy, at finite dielectric length, we observe that the discontinuity stops when $T^{*}\gtrsim T_{c}^{*}=0.183$, providing an estimate of the critical temperature. This value is in particular very close to the value obtained from BD simulations, $T_{c}^{*}=0.17$.
In order to compare our theoretical results with our BD simulations, we plot in Figure 3.(b) the evolution of the free ion fraction with the reduced Coulomb temperature for $\xi=13.8$ nm and for the perfect 2D case. To be consistent we use the same ionic density than BD simulations ($10^{16}$  atom.m-2). This density is below the tricritical point (for both system), and we indeed observe a transition at the critical temperature, between a phase with a low free ion fraction (fully paired for the perfect 2D case), and a phase with a large free ion fraction.
We finally note that for system with small $\xi$ such that liquid/wall with low dielectric contrast, or when the wall have a large Thomas-Fermi length Kavokine, Robin, and Bocquet (2022) (i.e. insulator or bad metal), the screening of the wall becomes dominant and the transition is not visible.
In Figure 3.(b), we plot the free ion fraction obtained in last subsections. Simulations and theories agrees very well qualitatively, in particular close to the transition. The discrepancies can be explained by the very different treatment of short range interactions in both system, yielding to different values of $K$ (defined in Equation (36)) and ionic radius $r_{0}$.
Figure 3: Theoretical predictions of the phase transition. (a) Evolution of free ion fraction with density below the transitions ($T^{*}=0.15$, in red) and above the transition ($T^{*}=0.35$, in blue), obtained from numerical solution of Equation (44). We use $\xi=13.8$ nm to match BD simulations. (b) Theoretical free ion fraction for the exact 2D Coulomb gas (red) and at finite dielectric length 13.8 nm (blue, comparable to BD simulations) and 27.6 nm (green, comparable to MD simulations), computed using Equation (44). The position of the transition temperature is shown by a cross. The result from respectively BD and MD simulations are plotted in blue and green, as well as the fit defined in Equation (40). In those cases the critical temperature obtained by the fit is indicated by a cross on the x axis.
II.4 Effect of the short-range interactions
II.4.1 Formation of ionic clusters

### v2

@@ I.3 Ionic current
where $\left\langle v_{+}\right\rangle$ and $\left\langle v_{-}\right\rangle$ are respectively the time-averaged velocity of the cations and the anions along the direction of the field, $N$ is the number of ions and $\mathcal{A}$ the 2D area spanned by the channel. We can deduce the conductivity through:
$\sigma(E)=\frac{j(E)}{E}.$ (7)
Which we will compare to the Ohmic conductance (defined for ions behaving like ideal tracer particles):
$\sigma_{0}=\frac{2N}{h\mathcal{A}}\frac{(Ze)^{2}D}{k_{B}T},$ (8)
with $D$ the diffusion coefficient of ions, assumed to be the same for anions and cations (in the case of a difference in diffusivity, $D$ should be replaced by the average of the coefficients of cations and anions). Parameters of BD simulations are chosen so that $D/k_{\text{B}}T$ is independent of temperature, so that we can use the same value of $\sigma_{0}$ to compare the conductivity of the system for different strengths of electrostatic interactions.
I.4 Ionic pairs and clusters
We observe in BD simulations that at low temperature, ions tend to form pairs. Since they are neutral and do not contribute to conduction, we use conductivity measurements as a way of quantifying the fraction of free ions $n_{f}$:
$n_{f}=\frac{\sigma(E)}{\sigma_{0}}.$ (9)
This definition of ionic pairing has the advantage of being the closest to what could be measured in experiments, in addition to being very convenient to compute. In fact, it effectively counts the fraction of charge carriers only if we can neglect higher-order correlations between ions (which tend to impede conduction compared to the Ohmic case), or equivalently if we assume the mobility of individual ions not to vary with the field. We make this assumption in the following, since we expect the non linearity of the system to be dominated by the variation of the fraction of charge carriers (second Wien effect), and not by the variation of the mobility (first Wien effectKaiser (2014); Kaiser et al. (2013), and Debye-Hückel correlations in generalBerthoumieux, Démery, and Maggs (2024)). This intuitive definition unfortunately fails if $E$ is too low, due to a poor signal-to-noise ratio. In SI, we report an alternative definition based on correlation functions, which is consistent with Eq. (9) but can be extended without problem to $E=0$.
In MD simulations, we observe the formation of large ionic clusters, in addition to ionic pairing. Some of these clusters bear a non-zero charge, and therefore participate to conduction. We therefore distinguish between two quantities: first, the fraction $n_{f}$ of free ions; and the fraction of charged clusters $n_{f}^{\text{cluster}}$ (which includes free ions – as they are charged clusters of size 1). In practice, we find that all clusters containing an even number of ions are electrically neutral, and odd-sized clusters have a charge ±$\pm 1$ (as clusters with multiple charge defects are highly unstable). From there, we can deduce that a measure of the fraction of charge carriers density is:
$n_{f}^{\text{cluster}}=\frac{\sum_{l\text{ odd}}N_{l}}{2N},$ (10)
with $N_{l}$ the number of clusters of size $l$, which we determine directly from trajectories through network analysis. The fraction of free ions is defined as:
$n_{f}=\frac{N_{1}}{2N}.$ (11)
This exact procedure used to compute $N_{l}$ from trajectories is described in the Supplementary Information.
II Pairing transition and electrostatic screening
@@
In this section, we quantify the formation of ionic pairs using both simulations and theoretical analysis. We start by predicting that the system undergoes a conductor/insulator phase transition reminiscent of the 2D Kosterlitz-Thouless transition. We analytically predict the system’s critical temperature, which is found to slightly deviate from the usual KT result due to the quasi-2D nature of confined electrolytes. We then compare those results to simulations. We find that indeed at low temperature, almost all ions assemble into pairs. We characterize this transition numerically, and show it is of infinite order as predicted by the KT analogy.
II.1 Theoretical analysis
II.1.1 Critical temperature and correlation length
It has been shownLevin (2002) that the 2D Coulomb gas, made of charged particles interacting through a potential $V(r)=-\frac{1}{T^{*}}\log r$, undergoes a pairing phase transition for $T^{*}<0.25$ that is equivalent to the Kosterlitez-Thouless transition. Here, we sketch the outline of such a derivation, while focusing on differences between the exact 2D Coulomb gas and our actual system – 2D confined electrolytes.
At the mean-field level, we can treat ion pairs as a separate, ideal chemical species, and write down the chemical equilibrium with free ions:
$n_{p}=1-n_{f}=\zeta n_{f}^{2}e^{2\beta\mu_{e}}.$ (12)
Here, $n_{p}$ is the fraction of ions that are part of a pair, $n_{f}$ is the free ion fraction, $\mu_{e}$ is the excess chemical potential of free ions, $\beta=1/k_{B}T$ and $\zeta$ is the internal partition function of a pair. In the 2D Coulomb gas, $\mu_{e}$ can be obtained through a Debye charging process, by computing the field created on a given ion by its surrounding Debye atmosphere:
$2\beta\mu_{e}=\frac{1}{T^{*}}\left[\frac{K_{0}(\kappa_{\text{D}}r_{0})}{\kappa_{\text{D}}r_{0}K_{1}(\kappa_{\text{D}}r_{0})}\right],$ (13)
where $r_{0}$ is a measure of the ionic size, $K_{0}$ and $K_{1}$ are modified Bessel functions of the second kind and
$\kappa_{\text{D}}=\sqrt{\frac{4\pi n_{f}c}{T^{*}}},$ (14)
is the inverse Debye length at salt concentration $c$ (expressed in atoms per surface area). For $T^{*}<T^{*}_{c}=0.25$, Eq. (12) becomes $1-n_{f}\propto n_{f}^{2-1/2T^{*}}$, with $02-1/2T^{*}<$ 0, and therefore admits no solution. This is the signature of the pairing transition. In addition, just above the transition, one can compute the correlation length through the Debye length (in the limit where the salt concentration is not too high Rotenberg, Bernard, and Hansen (2018)):
$\lambda_{\text{D}}=\sqrt{\frac{T^{*}}{4\pi n_{f}c}}\underset{T^{*}\to T^{*}_{c}}{\propto}e^{\frac{T^{*}/4}{T^{*}-T^{*}_{c}}}.$ (15)
The correlation length therefore depends in a non-algebraic manner on temperature, and in particular no critical exponent can be defined – this shows that the phase transition is of infinite order. Interestingly, the system possesses a critical line, meaning that there is a single value of the phase transition temperature regardless of salt concentration. This line ends at a given value of concentration, depending on the precise value of $\zeta$ and the exact criterion used to define ion pairs (for our choice of parameters, the transition occurs independently of concentration for $c<0.9\,$M).
In our case, however, ionic interactions are not exactly logarithmic: $V(r)\sim-\frac{1}{T^{*}}\log r/(r+\xi)$, see Eq. (3). Repeating the above argument, we find that:
$2\beta\mu_{e}=\frac{1}{T^{*}}\left[\frac{\xi}{\kappa_{\text{D}}r_{0}(r_{0}+\xi)}\frac{K_{0}(\kappa_{\text{D}}r_{0})-K_{0}(\kappa_{\text{D}}(r_{0}+\xi))}{K_{1}(\kappa_{\text{D}}r_{0})-K_{1}(\kappa_{\text{D}}(r_{0}+\xi))}\right].$ (16)
Then, Eq. (12) stops having a solution for
$T^{*}_{c}=-\frac{1}{4}\frac{\xi}{\xi+1}\min_{k}k\frac{\partial g_{\xi}}{\partial k}$ (17)
where we defined
$g_{\xi}(k)=\frac{K_{0}(k)-K_{0}(k(1+\xi/r_{0}))}{K_{1}(k)-K_{1}(k(1+\xi/r_{0}))}$ (18)
We notably find that $T^{*}_{c}\simeq 0.18$ for $\xi/r_{0}=14$ and we recover $T^{*}_{c}=1/4$ for $\xi\to\infty$.
Lastly, let us discuss the case where the dielectric length $\xi$ is short, such as for $\epsilon_{s}\sim\epsilon_{w}$. In that case, interactions are weakened as the electric field lines of ions leak into the walls. This situation is similar to confinement by conducting or metallic walls Kavokine, Robin, and Bocquet (2022); Rochester et al. (2013); Kondrat et al. (2011). While some degree of ionic pairing was reported in such context, it is a priori much weaker than for channels with high dielectric contrast such as here.
II.1.2 Discussion and scaling laws
In the previous paragraph we discussed the formation of ion pairs in confined electrolytes. These pairs, called Bjerrum pairs, are usually absent from bulk aqueous electrolytes; we thus discuss the impact of confinement below.
@@ II.1.2 Discussion and scaling laws
In any setting, the strength of the ionic interactions can be quantified by the Bjerrum lengthBjerrum (1926) $l_{B}$, defined as:
$\beta V(\ell_{\text{B}})=k_{\text{B}}T.$ (19)
It is the typical length at which electrostatic interactions become comparable to thermal agitation. For a bulk system, the Bjerrum length is then given by:
$\ell_{\text{B}}^{\text{bulk}}=\frac{e^{2}}{4\pi\epsilon_{0}\epsilon_{w}k_{B}T},$ (20)
while for a 2D confined electrolyte, where electrostatic interactions are quasi-logarithmic (see Eq. (3)), it reads:
$\ell_{\text{B}}^{\text{2D}}=\frac{\xi}{e^{T^{*}}-1}{\sim}\frac{\xi}{T^{*}}.$ (21)
@@
$\frac{\ell_{\text{B}}^{\text{2D}}}{\ell_{\text{B}}^{\text{bulk}}}\simeq\frac{\varepsilon_{w}}{\varepsilon_{s}}\gg 1.$ (22)
This notably shows how, under confinement, dielectric contrast between the walls and water greatly reinforces electostatic interactions. In bulk water, $\ell_{\text{B}}^{\text{bulk}}=0.7$nm, meaning that long-range interactions between ions are essentially negligible compared to thermal noise. In the slit this is not the case anymore, as the Bjerrum length is orders of magnitude larger: for our choice of parameters, we obtain $\ell_{\text{B}}=14\,$nm, far exceeding any other microscopic lengthscale. As we will discuss for several later points, 2D confined electrostatic interactions are nearly scale-invariant, owing to their quasi-logarithmic nature.
This fact allows us to understand the emergence of the phase transition using scaling laws. The excess chemical potential $\mu_{e}$ of free ions can be thought of as half the free energy cost of breaking a pair. Since the free cation and anion obtained by breaking a pair are on average distant by $\lambda_{\text{D}}$, we can assume that:
$2\beta\mu_{e}\sim\frac{1}{T^{*}}\log\lambda_{\text{D}}/r_{0},$ (23)
where $r_{0}$ is the ionic size. At low enough temperature, we expect that most ions are paired, so that $n_{p}=1-n_{f}\simeq 1$. From the chemical equilibrium (12), and using the definition of $\lambda_{\text{D}}$, we obtain:
$n_{f}\propto c^{\frac{1}{4T^{*}-1}}.$ (24)
This leads to a contradiction if $T^{*}<0.25$: for low enough salt concentration $c$, $n_{f}>1$, which is impossible – the chemical equilibrium is broken and there is a phase transition.
II.2 Comparison with simulations
II.2.1 Brownian dynamics simulations
In order to verify our theoretical results with our BD simulations, we plot in Figure 2.(a) the evolution of the free ion fraction with the reduced Coulomb temperature for $\xi=13.8$ nm. As we lower the temperature, the free ion fraction decrease, up to an almost fully paired system for a finite value of temperature. Quantitatively, one can define the crtitical temperature in BD simulations as the temperature for which the system’s correlation length diverges. In Fig. 2.(a), we show the evolution of the Debye length $\lambda_{\text{D}}$ with temperature, and we fit data from BD simulations with an ansatz inspired by Equation (15) (solid blue line):
$\lambda_{\text{D}}(T^{*})\propto\exp\left[\frac{T^{*}/4}{T^{*}-T^{*}_{c}}\right]$ (25)
We find $T_{c}^{*}=0.17$, which is very close to the theoretical value $T^{*}_{c}=0.18$. The comparison between theory and simulations notably confirms that the correlation length does diverges in a non-algebraic way close to the transition, validating the fact it is of infinite order. The comparison between the fraction of free ions in simulations and theory (see Equation (12)) shows similarly good agreement, as displayed on Figure 2.(d) (red symbols and solid line).
@@ II.2.1 Brownian dynamics simulations
Figure 2: Pairing transition in theory and simulations. (a) (Red points) Free ion fraction from BD simulation for various reduced Coulomb temperature. (Red line) Fit of the free ion fraction for an infinite order phase transition. (Blue points and line) Debye length obtained from the the free ion fraction using Equation (15). (b) Screenshot from MD simulations at equilibrium for $T^{*}=0.1$. We observe large ionic cluster (7 ions in the screenshot) that move together and than can have a net charge (-1 in the screenshot). (c) Schematic of the phase transition. At small reduced Coulomb temperature or equivalently high ionic interaction, the system is fully paired. Otherwise, there is some remaining free ions. (d) (Red points) Free ion fraction from BD simulations for various reduced Coulomb temperature. (Red line) Theoretical curve for $r_{0}=1.2$ nm and $\zeta=0.22$. (Green points) Free charge carrier density from MD simulations for various reduced Coulomb temperature. (Green line) Theoretical curve for $r_{0}=0.18$ nm and $\zeta=0.33$.
II.2.2 Molecular dynamics simulations and effect of the short-range interactions